# sample.bib

### v2

@@ 5.1 Comparisons with SOTAs
Figure 7: Visualization of embedding space learned under three encoder settings on the multi-organ segmentation (Synapse) test set. Distinct colors are used to differentiate embeddings from different encoders. Dimensions are reduced by t-SNE [van2008visualizing]. (a) Model-2 (Single encoder); (b) Model-3 (Dual encoders updated independently); (c) MOSformer (Dual encoders with a momentum update).
@@ 5.2 Ablation Study
Extensive ablation studies are conducted on the multi-organ segmentation (Synapse) and the automated cardiac diagnosis challenge (ACDC) datasets to verify the effectiveness of the momentum encoder and IF-Trans. DSC is selected as the default evaluation metric. Quantitative results are shown in Table 4. It should be noted that the baseline, Model-1, is a 2D-based model.
Importance of The Momentum Update. Two variants of MOSformer are employed in this experiment: i) Model-2: the encoder with a momentum update is removed, using a single encoder to extract features of target and neighborhood slices; ii) Model-3: the momentum encoder is replaced by a normal encoder and parameters of two encoders are updated independently via back-propagation. From quantitative results presented in Table 4, we can observe that these variants lead to decreased performance on both the Synapse dataset ($1.40\%$ and $0.70\%$ in DSC) and the ACDC dataset ($0.15\%$ and $0.09\%$ in DSC). The above results confirm the importance of the momentum update, designed to make slice features distinguishable and consistent. This design enables the model to distinguish target slices and fuse inter-slice information effectively.
Figure 8: Class activation maps of the gallbladder and the stomach category (from top to bottom) produced by Grad-CAM [selvaraju2017grad]. The class activation maps are generated from the last decoder layer.
@@
Furthermore, we also adopt t-SNE [van2008visualizing] to visualize the encoded embedding space learned from three encoder settings on the multi-organ segmentation (Synapse) test set. Model-2 employs a single encoder to process all input slices, where all slice features originate from the same feature space, as depicted in Fig. 7 (a). This setup poses challenges for the model in distinguishing individual slices and acquiring slice-specific information during inter-slice fusion. In contrast, the embedding space learned by dual encoders is distinguishable, as illustrated in Fig. 7 (b) and (c). It can also be discovered that incorporating the momentum update in dual encoders facilitates consistency among slice features, as shown in Fig. 7 (c), thereby further boosting segmentation performance.
Efficacy of The Inter-slice Fusion Transformer. We remove IF-Trans modules in MOSformer and obtain a baseline model (Model-1). Compared with the baseline model, models incorporating IF-Trans (Model-2 and MOSformer) offer substantial improvements on the Synapse ($1.81\%$ and $3.21\%$ DSC gains) and the ACDC dataset ($0.43\%$ and $0.58\%$ DSC gains). These results demonstrate inter-slice information is crucial in 3D medical image segmentation and the proposed IF-Trans can effectively fuse such information.
Additionally, we employ Grad-CAM [selvaraju2017grad] to visualize discriminative regions of the models, as depicted in Fig. 8. Compared with baseline Model-1, we can see that inter-slice information is beneficial, but Model-2 and Model-3 still tend to assign weights to irrelevant regions. Distinguishable and consistent inter-slice features within MOSformer can address the above issue, demonstrating enhanced precision in localizing organs of interest.
@@ 5.3 Hyperparameter Analysis
In this section, we conduct extensive analysis of several factors that correlate to segmentation performance of MOSformer. Default configurations of MOSformer are highlighted in gray.
@@
$/16$ 83.00 (-2.63) 21.54 (+8.14) 91.63 (-0.56) 1.14 (+0.06)
/8,/16/8,/16/ 8 , / 16 83.76 (-1.87) 20.73 (+7.33) 91.75 (-0.44) 1.08 (+0.00)
/4,/8,/16/4,/8,/16/ 4 , / 8 , / 16 84.52 (-1.11) 15.81 (+2.41) 91.94 (-0.25) 1.08 (+0.00)
/2,/4,/8,/16/2,/4,/8,/16/ 2 , / 4 , / 8 , / 16 85.63 13.40 92.19 1.08
Momentum Coefficient. The momentum coefficient, as described in Eq. (1), is an important hyperparameter in our model. We carry out detailed analysis on how $m$ affects the model performance, as shown in Fig. 9. Our empirical observations indicate a consistent decline in model performance with incremental increases in $m$. This suggests that maintaining feature consistency achieved through a relatively low momentum coefficient is advantageous. Specifically, a high momentum value (e.g., $m=0.9$) leads to a significant drop in segmentation performance, from $85.63\%$ to $83.49\%$ in DSC. At the extreme case of no momentum ($m=0$), the performance is nearly the poorest. These findings reinforce our motivation for extracting distinguishable and consistent slice features.
Neighborhood Slice Number. Since the proposed MOSformer is a 2.5D-based model, it requires neighborhood slices as additional inputs, as illustrated in Section 3. Thus, the number of neighborhood slices ($s$) is an important hyperparameter. Table 5 reports quantitative results for three different $s$ parameters. It can be observed that segmentation performance initially increases and then decreases with an increasing value of $s$. Evidently, information from inter-slice enables our model to perceive partial structures of 3D medical volumes. However, a peculiar phenomenon emerges: segmentation performance of the model with $s=2$ are worse than that with $s=1$. Similar observations have been reported in [zhang2022bridging]. One possible explanation is that the most valuable inter-slice information is derived from adjacent slices. Introducing non-adjacent slices may bring redundant information, which contributes negatively to model performance. Additionally, as $s$ increases, the computational costs of our model also escalate. Based on the above observations, $s=1$ is the most practical choice for our model.
@@
2D UNet [ronneberger2015u] [MICCAI’15] 17.26 30.74 0.67
TransUNet [chen2024transunet] [MedIA’24] 93.23 24.73 5.69
MISSformer [huang2023missformer] [TMI’23] 35.45 7.28 7.20
3D UNETR [hatamizadeh2022unetr] [WACV’22] 92.62 82.63 5.39
nnFormer [zhou2023nnformer] [TIP’23] 149.13 246.10 10.13
2.5D CAT-Net [hung2022cat] [TMI’22] 220.16 121.83 21.34
@@
Multi-scale Inter-slice Fusion. Multi-scale learning enables deep models to capture global spatial information and local contextual details. This conclusion has been supported by many studies [chen2024transunet][cao2022swin][you2022class]. In this paper, we further investigate multi-scale learning by incorporating inter-slice fusion. Table 6 presents results derived from four different inter-slice fusion configurations. Our default model achieves significant performance improvements, such as $1.11\%\sim 2.63\%$ gains in DSC on the multiorgan segmentation (Synapse) dataset and $0.25\%\sim 0.56\%$ gains in DSC on the automatic cardiac diagnosis (ACDC) dataset. With more scales of inter-slice information fused, MOSformer demonstrates an enhanced ability to comprehend global shapes and anatomical details within segmentation targets. This enhancement facilitates precise localization of semantic regions, resulting in higher DSC, and accurate classification of category boundaries, reflected in smaller HD95.
@@ 5.4 Model Complexity
Table 7 presents a comparison of five medical image segmentation models with MOSformer across various dimensions, including model parameters, floating-point operations per second (FLOPs), and the average time required for segmenting individual cases. MOSformer maintains a relatively small size (77.09 M) compared with 3D-based and 2.5D-based models. Furthermore, MOSformer exhibits an inference speed only half that of nnFormer [zhou2023nnformer], even surpassing 2D-based TransUNet [chen2024transunet] and MISSformer [huang2023missformer]. These results indicate MOSformer can achieve a favorable trade-off between model complexity and segmentation performance.
@@ 6 Conclusion
This paper proposes a MOmentum encoder-based inter-Slice fusion transformer (MOSformer) for stable and precise medical image segmentation. Dual encoders with a momentum update are able to guarantee both feature distinguishability and consistency, beneficial for inter-slice fusion. Besides, rich contexts can be captured via inter-slice self-attention in the IF-Trans module. The superior performance to state-of-the-arts on three benchmarks has demonstrated the MOSformer’s effectiveness and competitiveness. It will be extended to other downstream medical analysis tasks in our subsequent works.
@@ Acknowledgments
Funding
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFC2415100, in part by the National Natural Science Foundation of China under Grant 62222316, Grant 62373351, Grant 82327801, Grant 62073325, Grant 62303463, in part by the Chinese Academy of Sciences Project for Young Scientists in Basic Research under Grant No. YSBR-104, in part by the Beijing Natural Science Foundation under Grant F252068, Grant 4254107, in part by China Postdoctoral Science Foundation under Grant 2024M763535 and in part by CAMS Innovation Fund for Medical Sciences (ClFMS) under Grant 2023-I2M-C&T-B-017.
Author Contributions
D.X. Huang performed data curation, developed the methodology, conducted the experiments, and drafted the original manuscript. X.H. Zhou conceived the study and acquired funding, with additional funding support from M.J. Gui. X.L. Xie, S.Q. Liu, S.Y. Wang, and Z.Q. Feng reviewed and edited the manuscript. Z.G. Hou supervised the project. All authors approved the final version of the manuscript.
Conflicts of Interest
The authors declare that there is no conflict of interest regarding the publication of this article.

### v3

@@ 5.1 Comparisons with SOTAs
Figure 7: Visualization of embedding space learned under three encoder settings on the multi-organ segmentation (Synapse) test set (1,568 slices). Each point represents the feature of a slice. Distinct colors are used to differentiate embeddings from different encoders. Because Model-2 uses a single encoder for both target and neighborhood slices, their embeddings are identical. Therefore, only the blue points are shown. Dimensions are reduced by t-SNE [63]. (a) Model-2 (Single encoder); (b) Model-3 (Dual encoders updated independently); (c) MOSformer (Dual encoders with a momentum update).
@@ 5.2 Ablation Study
Extensive ablation studies are conducted on the multi-organ segmentation (Synapse), the automated cardiac diagnosis challenge (ACDC) and the abdominal organ segmentation (AMOS) datasets to verify the effectiveness of the momentum encoder and IF-Trans. DSC is selected as the default evaluation metric. Quantitative results are shown in Table 5. It should be noted that the baseline, Model-1, is a 2D-based model.
Importance of The Momentum Update. Two variants of MOSformer are employed in this experiment: i) Model-2: the encoder with a momentum update is removed, using a single encoder to extract features of target and neighborhood slices; ii) Model-3: the momentum encoder is replaced by a normal encoder and parameters of two encoders are updated independently via back-propagation. From quantitative results presented in Table 5, we can observe that these variants lead to decreased performance on the Synapse dataset ($1.40\%$ and $0.70\%$ in DSC), the ACDC dataset ($0.15\%$ and $0.09\%$ in DSC), and the AMOS dataset ($2.80\%$ and $1.55\%$ in DSC). The above results confirm the importance of the momentum update, designed to make slice features distinguishable and consistent. This design enables the model to distinguish target slices and fuse inter-slice information effectively.
Furthermore, we also adopt t-SNE [63] to visualize the encoded embedding space learned from three encoder settings on the multi-organ segmentation (Synapse) test set. Model-2 employs a single encoder to process both target and neighborhood slices. Consequently, target and neighborhood slice features originate from the same feature space, as depicted in Fig. 7 (a). This setup poses challenges for the model in distinguishing individual slices and acquiring slice-specific information during inter-slice fusion. In contrast, the embedding space learned by dual encoders is distinguishable, as illustrated in Fig. 7 (b) and (c). It can also be observed that incorporating the momentum update in dual encoders facilitates consistency among slice features, as shown in Fig. 7 (c), thereby further boosting segmentation performance.
Figure 8: Class activation maps of the gallbladder and the stomach categories (from top to bottom) produced by Grad-CAM [64]. The class activation maps are generated from the last decoder layer.
@@
Efficacy of The Inter-slice Fusion Transformer. We remove IF-Trans modules in MOSformer and obtain a baseline model (Model-1). Compared with the baseline model, models incorporating IF-Trans (Model-2 and MOSformer) offer substantial improvements on the Synapse ($1.81\%$ and $3.21\%$ DSC gains), the ACDC dataset ($0.43\%$ and $0.58\%$ DSC gains), and the AMOS dataset ($1.35\%$ and $4.15\%$ DSC gains). These results demonstrate that inter-slice information is crucial in 3D medical image segmentation and the proposed IF-Trans can effectively fuse such information.
Additionally, we employ Grad-CAM [64] to visualize discriminative regions of the models, as depicted in Fig. 8. Compared with baseline Model-1, we can see that inter-slice information is beneficial, but Model-2 and Model-3 still tend to assign weights to irrelevant regions. Distinguishable and consistent inter-slice features within MOSformer can address the above issue, demonstrating enhanced precision in localizing organs of interest.
@@ 5.3 Hyperparameter Analysis
In this section, we conduct extensive analysis of several factors that correlate with segmentation performance of MOSformer. Default configurations of MOSformer are highlighted in gray.
Momentum Coefficient. The momentum coefficient, as described in Eq. (1), is an important hyperparameter in our model. We carry out detailed analysis on how $m$ affects the model performance, as shown in Fig. 9. Our empirical observations indicate a consistent decline in model performance with incremental increases in $m$. This suggests that maintaining feature consistency achieved through a relatively low momentum coefficient is advantageous. Specifically, a high momentum value (e.g., $m=0.9$) leads to a significant drop in segmentation performance, from $85.63\%$ to $83.49\%$ in DSC. In the extreme case of no momentum ($m=0$), the performance is nearly the worst. These findings reinforce our motivation for extracting distinguishable and consistent slice features.
Neighborhood Slice Number. Since the proposed MOSformer is a 2.5D-based model, it requires neighborhood slices as additional inputs, as illustrated in Section 3. Thus, the number of neighborhood slices ($s$) is an important hyperparameter. Table 6 reports quantitative results for three different $s$ parameters. It can be observed that segmentation performance initially increases and then decreases with an increasing value of $s$. Evidently, information from inter-slice enables our model to perceive partial structures of 3D medical volumes. However, a peculiar phenomenon emerges: segmentation performance of the model with $s=2$ is worse than that with $s=1$. Similar observations have been reported in [18]. One possible explanation is that the most valuable inter-slice information is derived from adjacent slices. Introducing non-adjacent slices may bring redundant information, which contributes negatively to model performance. Additionally, as $s$ increases, the computational costs of our model also escalate. Based on the above observations, $s=1$ is the most practical choice for our model.
@@
/16/16 83.00 (-2.63) 21.54 (+8.14) 91.63 (-0.56) 1.14 (+0.06)
/8,/16/8,/16 83.76 (-1.87) 20.73 (+7.33) 91.75 (-0.44) 1.08 (+0.00)
/4,/8,/16/4,/8,/16 84.52 (-1.11) 15.81 (+2.41) 91.94 (-0.25) 1.08 (+0.00)
/2,/4,/8,/16/2,/4,/8,/16 85.63 13.40 92.19 1.08
@@
2D UNet [12] [MICCAI’15] 17.26 30.74 0.67
TransUNet [23] [MedIA’24] 93.23 24.73 5.69
MISSformer [43] [TMI’23] 35.45 7.28 7.20
3D UNETR [44] [WACV’22] 92.62 82.63 5.39
nnFormer [46] [TIP’23] 149.13 246.10 10.13
2.5D CAT-Net [28] [TMI’22] 220.16 121.83 21.34
@@
Multi-scale Inter-slice Fusion. Multi-scale learning enables deep models to capture global spatial information and local contextual details. This conclusion has been supported by many studies [23][22][42]. In this paper, we further investigate multi-scale learning by incorporating inter-slice fusion. Table 7 presents results derived from four different inter-slice fusion configurations. Our default model achieves significant performance improvements, such as $1.11\%\sim+2.63\%$ gains in DSC on the multiorgan segmentation (Synapse) dataset and $0.25\%\sim+0.56\%$ gains in DSC on the automatic cardiac diagnosis (ACDC) dataset. With more scales of inter-slice information fused, MOSformer demonstrates an enhanced ability to comprehend global shapes and anatomical details within segmentation targets. This enhancement facilitates precise localization of semantic regions, resulting in higher DSC, and accurate classification of category boundaries, reflected in smaller HD95.
@@ 5.4 Model Complexity
Table 8 presents a comparison of five medical image segmentation models with MOSformer across various dimensions, including model parameters, floating-point operations per second (FLOPs), and the average time required for segmenting individual cases. MOSformer maintains a relatively small size (77.09 M) compared with 3D-based and 2.5D-based models. Furthermore, MOSformer exhibits an inference speed only half that of nnFormer [46], even surpassing 2D-based TransUNet [23] and MISSformer [43]. These results indicate MOSformer can achieve a favorable trade-off between model complexity and segmentation performance.
@@ 6 Conclusion
This study proposes a MOmentum encoder-based inter-Slice fusion transformer (MOSformer) for stable and precise medical image segmentation. Dual encoders with a momentum update are able to guarantee both feature distinguishability and consistency, beneficial for inter-slice fusion. Besides, rich contexts can be captured via inter-slice self-attention in the IF-Trans module. The superior performance to state-of-the-art methods on three benchmarks has demonstrated MOSformer’s effectiveness and competitiveness. It will be extended to other downstream medical analysis tasks in our subsequent works.
[…]
@@ Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFC2415100, in part by the National Natural Science Foundation of China under Grant 62222316, Grant 62373351, Grant 82327801, Grant 62073325, Grant 62303463, in part by the Chinese Academy of Sciences Project for Young Scientists in Basic Research under Grant No. YSBR-104, in part by the Beijing Natural Science Foundation under Grant F252068, Grant 4254107, in part by Beijing Nova Program under Grant 20250484813, in part by China Postdoctoral Science Foundation under Grant 2024M763535, in part by the Postdoctoral Fellowship Program of CPSF under Grant GZC20251170 and in part by CAMS Innovation Fund for Medical Sciences (ClFMS) under Grant 2023-I2M-C&T-B-017.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.